# A Fisher-Rao gradient flow for entropic mean-field min-max games

### v1

@@ Abstract.
1. Introduction
The rapid progress of machine learning (ML) techniques such as Generative Adversarial Networks (GANs) [6], adversarial learning [16], multi-agent reinforcement learning [27] has propelled a surge of interest in the study of optimization problems on the space of probability measures in recent years. Particularly noteworthy are the numerous works, e.g., [7, 4, 26, 12, 21, 8], illustrating how training GANs and addressing adversarial robustness can be cast as min-max games over probability measures. In this setting, understanding the time evolution of the players’ initial strategies to the equilibrium of the game leverages the use of gradient flows on the space of probability measures. A discussion about the applications of gradient flows in optimization and sampling can be found in a recent survey [22].
The Fisher-Rao (FR) gradient flow has been recently studied in the context of mean-field optimization problems [11], accelerating Langevin-based sampling from multi-modal distributions [13, 14], and training shallow neural networks in the mean-field regime [19]. The present paper aims to extend these results and focuses on the continuous-time convergence of the Fisher-Rao (FR) gradient flow to the unique mixed Nash equilibrium of an entropy-regularized min-max game.
1.1. Notation and setup
@@
(1) $\min_{\nu\in\mathcal{P}\left(\mathcal{X}\right)}\max_{\mu\in\mathcal{P}\left(\mathcal{Y}\right)}V^{\sigma}(\nu,\mu),\text{ with }V^{\sigma}(\nu,\mu)\coloneqq F(\nu,\mu)+\frac{\sigma^{2}}{2}\left(\operatorname{D_{KL}}(\nu|\pi)-\operatorname{D_{KL}}(\mu|\rho)\right).$
In order to solve (1), one typically aims to identify mixed Nash equilibria (MNEs) [24, 17], characterized by pairs of measures $(\nu_{\sigma}^{*},\mu_{\sigma}^{*})\in\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{Y})$ that satisfy
(2) $V^{\sigma}(\nu_{\sigma}^{*},\mu)\leq V^{\sigma}(\nu_{\sigma}^{*},\mu_{\sigma}^{*})\leq V^{\sigma}(\nu,\mu_{\sigma}^{*}),\quad\text{for all }(\nu,\mu)\in\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{Y}).$
It is important to highlight that when $F$ is bilinear and $\sigma=0$, i.e.  $V^{0}(\nu,\mu)=\int_{\mathcal{Y}}\int_{\mathcal{X}}f(x,y)\nu(\mathrm{d}x)\mu(\mathrm{d}y),$ for some $f:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$, measures characterized by (2) represent MNEs in the classical sense of a two-player zero-sum game. Results concerned with the existence and uniqueness of an MNE for (1) are presented in Appendix A in [9]. Therein, it is also proved that $V^{\sigma}$ $\Gamma$-converges to $F$ as $\sigma\downarrow 0,$ under mild assumptions on $F,\pi$ and $\rho.$
In optimization, the monotonic decrease of the objective function along the gradient flow is key to proving convergence. However, for min-max games the monotonic decrease no longer holds due to the conflicting actions of the players. Hence, a suitable Lyapunov function is needed. A common choice is the so-called Nikaidò-Isoda (NI) error [18], which, for all $(\nu,\mu)\in\mathcal{P}\left(\mathcal{X}\right)\times\mathcal{P}\left(\mathcal{Y}\right)$, can be defined as
@@
1.2. Fisher-Rao (mean-field birth-death) gradient flow
In this work, we are mainly focusing on the the dynamical formulation of the Fisher-Rao metric (see e.g. [5, 11]). For any $\lambda_{0}$, $\lambda_{1}\in\mathcal{P}_{\text{ac}}(\mathcal{M})$, with $\mathcal{M}\subseteq\mathbb{R}^{d},$ the variational representation of the FR distance is given by
(3) $\operatorname{FR}^{2}(\lambda_{0},\lambda_{1})\coloneqq\inf\Bigg{\{}\int_{0}^{1}\int_{\mathcal{M}}\left|r_{s}(x)-\int_{\mathcal{M}}r_{s}(y)\lambda_{s}(\mathrm{d}y)\right|^{2}\lambda_{s}(\mathrm{d}x)\mathrm{d}s:\\ \partial_{t}\lambda_{t}=\lambda_{t}\left(r_{t}-\int_{\mathcal{M}}r_{t}(y)\lambda_{t}(\mathrm{d}y)\right)\Bigg{\}},$
@@ 1.2. Fisher-Rao (mean-field birth-death) gradient flow
(4) $\partial_{t}\lambda_{t}=\lambda_{t}\left(r_{t}-\int_{\mathcal{M}}r_{t}(y)\lambda_{t}(\mathrm{d}y)\right)$
@@
Inspired by [11], we consider the Fisher-Rao (mean-field birth-death) gradient flow on the space $\left(\mathcal{P}_{\text{ac}}(\mathcal{X})\times\mathcal{P}_{\text{ac}}(\mathcal{Y}),\text{FR}\right)$ in the setting of (1). As opposed to the mean-field Best Response dynamics studied in [9], which is another flow that can be used to solve (1), and which relies on introducing a fixed point perspective on min-max games, the FR dynamics utilize a gradient flow $(\nu_{t},\mu_{t})_{t\geq 0}$ in the Fisher-Rao geometry. As a first attempt at defining a Fisher-Rao gradient flow for solving (1), consider
(5) $\begin{cases}\partial_{t}\nu_{t}(x)=-\frac{\delta V^{\sigma}}{\delta\nu}(\nu_{t},\mu_{t},x)\nu_{t}(x),\\ \partial_{t}\mu_{t}(y)=\frac{\delta V^{\sigma}}{\delta\mu}(\nu_{t},\mu_{t},y)\mu_{t}(y),\end{cases}$
@@
since the flat derivatives of $V^{\sigma}$ are uniquely defined up to an additive shift (see Definition B.1 in Section B). Thus, the total mass $1$ of probability measures is still preserved along the gradient flow.
1.2.1. Sketch of convergence proof for the FR gradient flow
For the sake of presenting an intuitive heuristic argument, we ignore here for now that the flat derivatives $(\nu,\mu,x)\mapsto\frac{\delta V^{\sigma}}{\delta\nu}(\nu,\mu,x)$ and $(\nu,\mu,y)\mapsto\frac{\delta V^{\sigma}}{\delta\mu}(\nu,\mu,y)$ may not exist for the $V^{\sigma}$ defined in (1), due to the relative entropy term $\operatorname{D_{KL}}$ being only lower semicontinuous (for this reason, in our analysis in Section 2, we will replace these two derivatives with appropriately defined auxiliary functions $a$ and $b$). Nevertheless, we will now demonstrate that choosing the flow $(\nu_{t},\mu_{t})_{t\geq 0}$ as in (5), makes the function
(6) $t\mapsto\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{t})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{t})$
@@ 1.2.1. Sketch of convergence proof for the FR gradient flow
(7) $\frac{\mathrm{d}}{\mathrm{d}t}\left(\operatorname{D_{KL}}(\nu|\nu_{t})+\operatorname{D_{KL}}(\mu|\mu_{t})\right)\leq V^{\sigma}(\nu,\mu_{t})-V^{\sigma}(\nu_{t},\mu_{t})+V^{\sigma}(\nu_{t},\mu_{t})-V^{\sigma}(\nu_{t},\mu)\\ -\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\nu|\nu_{t})-\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\mu|\mu_{t}).$
@@
$\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{t})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{t})\leq e^{-\frac{\sigma^{2}}{2}t}\left(\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{0})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{0})\right).$
@@
Since $\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu)+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu)\geq 0$ with equality if and only if $\nu=\nu_{\sigma}^{*}$ and $\mu=\mu_{\sigma}^{*}$, it follows that the unique MNE of (1) is achieved with exponential rate $e^{-\frac{\sigma^{2}}{2}t}.$ To prove convergence in terms of the NI error, first observe that since
$\frac{\sigma^{2}}{2}\left(\operatorname{D_{KL}}(\nu|\nu_{t})+\operatorname{D_{KL}}(\mu|\mu_{t})\right)\geq 0,$
for all $(\nu,\mu),$ (7) becomes
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\operatorname{D_{KL}}(\nu|\nu_{t})+\operatorname{D_{KL}}(\mu|\mu_{t})\right)\leq V^{\sigma}(\nu,\mu_{t})-V^{\sigma}(\nu_{t},\mu).$
We integrate this inequality from 00 to $t>0,$ divide by $t$ and apply Jensen’s inequality to $\nu\mapsto V^{\sigma}(\nu,\mu)$ and $\mu\mapsto V^{\sigma}(\nu,\mu),$ respectively. Lastly, maximizing over $(\nu,\mu)$ gives
$\operatorname{NI}\left(\frac{1}{t}\int_{0}^{t}\nu_{s}\mathrm{d}s,\frac{1}{t}\int_{0}^{t}\mu_{s}\mathrm{d}s\right)\leq\frac{1}{t}\left(\max_{\nu}\operatorname{D_{KL}}(\nu|\nu_{0})+\max_{\mu}\operatorname{D_{KL}}(\mu|\mu_{0})\right).$
1.3. Our contribution
@@
We prove the existence of the FR gradient flow $(\nu_{t},\mu_{t})_{t\geq 0}$ and show that it converges with rate $\mathcal{O}\left(e^{-\frac{\sigma^{2}}{2}t}\right)$ to the unique MNE of (1) with respect to the Lyapunov function $t\mapsto\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{t})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{t}).$ We also show that $\operatorname{NI}\left(\frac{1}{t}\int_{0}^{t}\nu_{s}\mathrm{d}s,\frac{1}{t}\int_{0}^{t}\mu_{s}\mathrm{d}s\right)$ converges to zero with rate $\mathcal{O}\left(\frac{1}{t}\right).$
1.4. Related works

### v2

@@ Abstract
1 Introduction
The rapid progress of machine learning (ML) techniques such as Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), adversarial learning (Madry et al., 2018), multi-agent reinforcement learning (Zhang et al., 2021) has propelled a surge of interest in the study of optimization problems on the space of probability measures in recent years. Particularly noteworthy are the numerous works, e.g., (Hsieh et al., 2019; Domingo-Enrich et al., 2020; Wang & Chizat, 2023; Lu, 2023; Trillos & Trillos, 2023; Kim et al., 2024), illustrating how training GANs and addressing adversarial robustness can be cast as min-max games over probability measures. In this setting, understanding the time evolution of the players’ initial strategies to the equilibrium of the game leverages the use of gradient flows on the space of probability measures. A discussion about the applications of gradient flows in optimization and sampling can be found in a recent survey (Trillos et al., 2023).
The Fisher-Rao (FR) gradient flow has been recently studied in the context of mean-field optimization problems (Liu et al., 2023), accelerating Langevin-based sampling from multi-modal distributions (Lu et al., 2019; 2023), and training shallow neural networks in the mean-field regime (Rotskoff et al., 2019). The motivation for employing FR dynamics in sampling from multi-modal distributions is their ability to globally transport the mass of a probability density between modes without traversing low-probability regions (Lu et al., 2019). In the context of training neural networks, the benefit of using FR dynamics is similar in that they have the potential capability to avoid local minima of the loss function (Rotskoff et al., 2019).
The present paper aims to extend these results and focuses on the continuous-time convergence of the Fisher-Rao (FR) gradient flow to the unique mixed Nash equilibrium of an entropy-regularized min-max game.
1.1 Notation and setup
@@
$\min_{\nu\in\mathcal{P}\left(\mathcal{X}\right)}\max_{\mu\in\mathcal{P}\left(\mathcal{Y}\right)}V^{\sigma}(\nu,\mu),\text{ with }V^{\sigma}(\nu,\mu)\coloneqq F(\nu,\mu)+\frac{\sigma^{2}}{2}\left(\operatorname{D_{KL}}(\nu|\pi)-\operatorname{D_{KL}}(\mu|\rho)\right).$ (1)
In order to solve (1), one typically aims to identify mixed Nash equilibria (MNEs) (von Neumann et al., 1944; Nash, 1951), characterized by pairs of measures $(\nu_{\sigma}^{*},\mu_{\sigma}^{*})\in\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{Y})$ that satisfy
$V^{\sigma}(\nu_{\sigma}^{*},\mu)\leq V^{\sigma}(\nu_{\sigma}^{*},\mu_{\sigma}^{*})\leq V^{\sigma}(\nu,\mu_{\sigma}^{*}),\quad\text{for all }(\nu,\mu)\in\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{Y}).$ (2)
It is important to highlight that when $F$ is bilinear and $\sigma=0$, i.e., $V^{0}(\nu,\mu)=\int_{\mathcal{Y}}\int_{\mathcal{X}}f(x,y)\nu(\mathrm{d}x)\mu(\mathrm{d}y),$ for some measurable function $f:\mathcal{X}\times\mathcal{Y}\to\mathbb{R},$ measures characterized by (2) represent MNEs in the classical sense of a two-player zero-sum game. Results concerned with the existence and uniqueness of an MNE for (1) are presented in Appendix A in Lascu et al. (2023) (alternatively, see Appendix C.2, C.3 in Domingo-Enrich et al. (2020) for the case when $F$ is bilinear). It is also proved in Lascu et al. (2023) that $V^{\sigma}$ $\Gamma$-converges to $F$ as $\sigma\downarrow 0,$ under mild assumptions on $F,\pi$ and $\rho.$
In optimization, the monotonic decrease of the objective function along the gradient flow is key to proving convergence. However, for min-max games the monotonic decrease no longer holds due to the conflicting actions of the players. Hence, a suitable Lyapunov function is needed. A common choice is the so-called Nikaidò-Isoda (NI) error (Nikaidô & Isoda, 1955), which, for all $(\nu,\mu)\in\mathcal{P}\left(\mathcal{X}\right)\times\mathcal{P}\left(\mathcal{Y}\right)$, can be defined as
@@
1.2 Fisher-Rao (mean-field birth-death) gradient flow
The Fisher-Rao metric was introduced by Rao (1992) via the Fisher information matrix, and since then has been extensively studied in the context of information geometry (Amari, 2016; Ay et al., 2017). In this work, we are mainly focusing on the the dynamical formulation of the Fisher-Rao metric (see e.g. Section 2.2 in Gallouët & Monsaingeon (2017), Appendix C.2.1 in Yan et al. (2023) and Section 3.3 in Kondratyev & Vorotnikov (2019)). For any $\lambda_{0}$, $\lambda_{1}\in\mathcal{P}_{\text{ac}}(\mathcal{M})$, with $\mathcal{M}\subseteq\mathbb{R}^{d},$ the variational representation of the FR distance is given by
$\operatorname{FR}^{2}(\lambda_{0},\lambda_{1})\coloneqq\inf\Bigg{\{}\int_{0}^{1}\int_{\mathcal{M}}\left|r_{s}(x)-\int_{\mathcal{M}}r_{s}(y)\lambda_{s}(\mathrm{d}y)\right|^{2}\lambda_{s}(\mathrm{d}x)\mathrm{d}s:\partial_{t}\lambda_{t}=\lambda_{t}\left(r_{t}-\int_{\mathcal{M}}r_{t}(y)\lambda_{t}(\mathrm{d}y)\right)\Bigg{\}},$ (3)
@@ 1.2 Fisher-Rao (mean-field birth-death) gradient flow
$\partial_{t}\lambda_{t}=\lambda_{t}\left(r_{t}-\int_{\mathcal{M}}r_{t}(y)\lambda_{t}(\mathrm{d}y)\right)$ (4)
@@
Inspired by Liu et al. (2023), we consider the Fisher-Rao (mean-field birth-death) gradient flow on the space $\left(\mathcal{P}_{\text{ac}}(\mathcal{X})\times\mathcal{P}_{\text{ac}}(\mathcal{Y}),\text{FR}\right)$ in the setting of (1). As opposed to the mean-field Best Response dynamics studied in Lascu et al. (2023), which is another flow that can be used to solve (1), and which relies on introducing a fixed point perspective on min-max games, the FR dynamics utilize a gradient flow $(\nu_{t},\mu_{t})_{t\geq 0}$ in the Fisher-Rao geometry. As a first attempt at defining a Fisher-Rao gradient flow for solving (1), consider
$\begin{cases}\partial_{t}\nu_{t}(x)=-\frac{\delta V^{\sigma}}{\delta\nu}(\nu_{t},\mu_{t},x)\nu_{t}(x),\\ \partial_{t}\mu_{t}(y)=\frac{\delta V^{\sigma}}{\delta\mu}(\nu_{t},\mu_{t},y)\mu_{t}(y),\end{cases}$ (5)
@@
since the flat derivatives of $V^{\sigma}$ are uniquely defined up to an additive shift (see Definition B.1 in Section B). Thus, the total mass $1$ of probability measures is still preserved along the gradient flow. A formal derivation of the Fisher-Rao gradient flow can be found in Appendix C.
Remark 1.1.
Suppose $V^{0}$ is a bilinear payoff function and any mixed strategies $(\nu,\mu)$ are supported on finite sets of $m$ and $n$ pure strategies $\mathcal{X}\coloneqq\{x_{1},x_{2},\cdots,x_{m}\}$ and $\mathcal{Y}\coloneqq\{y_{1},y_{2},\cdots,y_{n}\},$ respectively. Then $\nu\in\mathcal{P}(\mathcal{X})$ and $\mu\in\mathcal{P}(\mathcal{Y})$ can be approximated by the empirical measures $\nu^{m}\coloneqq\frac{1}{m}\sum_{i=1}^{m}\delta_{x_{i}}$ and $\mu^{n}\coloneqq\frac{1}{n}\sum_{j=1}^{n}\delta_{y_{j}},$ respectively, and hence the Fisher-Rao gradient flow (5) retrieves the replicator dynamics studied in evolutionary game theory (Cressman & Tao, 2014; Hofbauer & Sigmund, 1998), see also (Abe et al., 2022) for a related replicator-mutant dynamics.
1.2.1 Sketch of convergence proof for the FR gradient flow
For the sake of presenting an intuitive heuristic argument, we ignore here for now that the flat derivatives $(\nu,\mu,x)\mapsto\frac{\delta V^{\sigma}}{\delta\nu}(\nu,\mu,x)$ and $(\nu,\mu,y)\mapsto\frac{\delta V^{\sigma}}{\delta\mu}(\nu,\mu,y)$ may not exist for the $V^{\sigma}$ defined in (1), due to the relative entropy term $\operatorname{D_{KL}}$ being only lower semicontinuous with respect to the weak convergence topology (for this reason, in our analysis in Section 2, we will replace these two derivatives with appropriately defined auxiliary functions $a$ and $b$). Nevertheless, we will now demonstrate that choosing the flow $(\nu_{t},\mu_{t})_{t\geq 0}$ as in (5), makes the function
$t\mapsto\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{t})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{t})$ (6)
@@ 1.2.1 Sketch of convergence proof for the FR gradient flow
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\operatorname{D_{KL}}(\nu|\nu_{t})+\operatorname{D_{KL}}(\mu|\mu_{t})\right)\leq V^{\sigma}(\nu,\mu_{t})-V^{\sigma}(\nu_{t},\mu_{t})+V^{\sigma}(\nu_{t},\mu_{t})-V^{\sigma}(\nu_{t},\mu)\\ -\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\nu|\nu_{t})-\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\mu|\mu_{t}).$ (7)
@@
$\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{t})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{t})\leq e^{-\frac{\sigma^{2}}{2}t}\left(\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{0})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{0})\right).$ (8)
@@
Since $\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu)+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu)\geq 0$ with equality if and only if $\nu=\nu_{\sigma}^{*}$ and $\mu=\mu_{\sigma}^{*}$, it follows that the unique MNE of (1) is achieved with exponential rate $e^{-\frac{\sigma^{2}}{2}t}.$
To prove convergence in terms of the NI error, first observe that, for any $(\nu,\mu)\in\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{Y}),$ we can write
$V^{\sigma}(\nu_{t},\mu)-V^{\sigma}(\nu,\mu_{t})=V^{\sigma}(\nu_{t},\mu)-V^{\sigma}(\nu_{\sigma}^{*},\mu)+V^{\sigma}(\nu_{\sigma}^{*},\mu)-V^{\sigma}(\nu,\mu_{\sigma}^{*})+V^{\sigma}(\nu,\mu_{\sigma}^{*})-V^{\sigma}(\nu,\mu_{t}).$
By Theorem 2.2 and Assumption 2, there exist $C_{1,\sigma},C_{2,\sigma}>0$ depending on $\sigma$ such that
$\left|\frac{\delta V^{\sigma}}{\delta\nu}(\nu_{t},\mu,x)\right|\leq C_{1,\sigma},\quad\left|\frac{\delta V^{\sigma}}{\delta\mu}(\nu,\mu_{t},y)\right|\leq C_{2,\sigma},$
for all $(\nu,\mu)\in\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{Y}),$ and all $(x,y)\in\mathcal{X}\times\mathcal{Y}.$ Then, by Lemma 3.1 and Pinsker’s inequality, we can show that
$V^{\sigma}(\nu_{t},\mu)-V^{\sigma}(\nu_{\sigma}^{*},\mu)+V^{\sigma}(\nu,\mu_{\sigma}^{*})-V^{\sigma}(\nu,\mu_{t})\leq 2C_{\sigma}\sqrt{\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{t})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{t})},$
where $C_{\sigma}\coloneqq\max\{C_{1,\sigma},C_{2,\sigma}\}.$ Hence, by (8), we obtain
$V^{\sigma}(\nu_{t},\mu)-V^{\sigma}(\nu,\mu_{t})\leq 2C_{\sigma}e^{-\frac{\sigma^{2}}{4}t}\sqrt{\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{0})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{0})}+V^{\sigma}(\nu_{\sigma}^{*},\mu)-V^{\sigma}(\nu,\mu_{\sigma}^{*}).$ (9)
Since $(\nu_{\sigma}^{*},\mu_{\sigma}^{*})$ is the MNE of (1), we have
$\max_{\mu}V^{\sigma}(\nu_{\sigma}^{*},\mu)=\min_{\nu}V^{\sigma}(\nu,\mu_{\sigma}^{*})=V^{\sigma}(\nu_{\sigma}^{*},\mu_{\sigma}^{*}).$
Therefore, maximizing over $(\nu,\mu)\in\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{Y})$ in (9) gives
$\operatorname{NI}(\nu_{t},\mu_{t})\leq 2C_{\sigma}e^{-\frac{\sigma^{2}}{4}t}\sqrt{\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{0})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{0})}.$
Finally, observe that, by (18), we have
$\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{0})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{0})<\infty.$
1.3 Our contribution
@@
We prove the existence of the FR gradient flow $(\nu_{t},\mu_{t})_{t\geq 0}$ and show that it converges with rates $\mathcal{O}\left(e^{-\frac{\sigma^{2}}{2}t}\right)$ and $\mathcal{O}\left(e^{-\frac{\sigma^{2}}{4}t}\right)$ to the unique MNE of (1) with respect to the Lyapunov functions $t\mapsto\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{t})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{t})$ and $t\mapsto\operatorname{NI}\left(\nu_{t},\mu_{t}\right).$
1.4 Related works